### v2

# Bruno Bauwens 11 $0$-0002-6138-0591    Maria Marchenko 22

Abstract
The even online Kolmogorov complexity of a string $x=x_{1}x_{2}\cdots x_{n}$ is the minimal length of a program that for all $i\leq n/2$, on input $x_{1}x_{3}\cdots x_{2i-1}$ outputs $x_{2i}$. The odd complexity is defined similarly. The sum of the odd and even complexities is called the dialogue complexity.
In [Bauwens, 2014] it is proven that for all $n$, there exist $n$-bit $x$ for which the dialogue complexity exceeds the Kolmogorov complexity by $n\log\frac{4}{3}+O(\log n)$. Let $\operatorname{C}^{s}(x)$ denote the Kolmogorov complexity with space bound $s$. Here, we prove that the space-bounded dialogue complexity with bound $s+6n+O(1)$ is at most $\operatorname{C}^{s}(x)+O(\log(sn))$, where $n=|x|$.
Keywords:
Online Kolmogorov complexity Symmetry of information Space-bounded computation
1 Introduction
The Kolmogorov complexity of a bitstring $x$ is the minimal length of a program that prints $x$. It was first used by Solomonoff to define an ideal algorithm for general sequence prediction [19, 20, 21]. However, in real life situations, prediction often exploits side information. For example, one may aim to predict whether certain stocks go up or down relative to macro economic indicators. This corresponds to a complexity notion where one alternates between receiving information and generating output.
The online Kolmogorov complexity is a generalization of conditional complexity, introduced in [7]. Given two lists of strings $x_{1},\ldots,x_{k}$ and $y_{1},\ldots,y_{k}$, it is defined by the minimal length of a program that for all $j\leq k$, on input $y_{1},\ldots,y_{j}$ prints $x_{j}$. This notion has important applications in the foundations of reinforcement learning [11, 12].111 An agent and an environment exchange messages. The AIXI-agent defined in [11, 12] models the environment using online a priori probability, which is a probabilistic version of complexity. The online coding theorem provides a close relation between the probabilistic and program complexity versions. This theorem is essential to prove the main result.
A special case is the even complexity $\operatorname{C}_{\textnormal{ev}}(x)$, which is the task of predicting each next even bit in $x$ given previous odd bits. The odd complexity $\operatorname{C}_{\textnormal{odd}}(x)$ is defined similarly. A time-bounded variant of this complexity is used in a recent combinatorial characterization of 2 cryptographic conjectures: the existence of public key cryptography and one way functions [1].
Kolmogorov complexity satisfies an additivity property: the information in a pair $(x,y)$ of bitstrings equals the information in $x$ and the information in $y$ given $x$, thus
$\operatorname{C}(x,y)=\operatorname{C}(x)+\operatorname{C}(y|x)+O(\log\operatorname{C}(x,y)).$ (*)
Note that for Shannon entropy, the chain rule is exact $H(X,Y)=H(X)+H(Y|X)$. The additional non constant big-O term is (among other things) a consequence of the uncomputability of the complexity function [6, 8]. Remarkably, such a result does not hold for online complexity: the sum of odd and even complexities can be much larger than the complexity of a string. Again, this is a consequence of the uncomputability of complexity.
Theorem 1.1 ([4, 16])
There exists a sequence $\omega_{1}\omega_{2}\cdots$ such that for all $n$,
$(\operatorname{C}_{\textnormal{ev}}+\operatorname{C}_{\textnormal{odd}})(\omega_{1}\cdots\omega_{n})\geq n\log\tfrac{4}{3}+\operatorname{C}(\omega_{1}\cdots\omega_{n}).$
There is a close relation to Muchnik’s paradox [16] in algorithmic randomness: if an infinite sequence is not random, one might neither blame the even nor the odd bits. Similar results for computable randomness and Hausdorff dimensions are given in [2, 3].
In this paper, we study additivity properties for space bounded variants. Note that there is no hope to prove a time-bounded version of this result, because it would contradict the existence of one-way functions [10, 14, 15].
The Kolmogorov complexity $\operatorname{C}^{s}(x|y)$ with space-bound $s$, is the minimal length of a program that prints $x$ on input $y$ and uses at most $s$ cells of memory [5, 9], (see section 2 for the full definition). It is well known that this complexity satisfies an additivity property. Note that the equality (*1) can be viewed as two inequalities. With space-bounded complexity, the smaller side of each inequality needs a larger space-bound. The easy direction is
$\operatorname{C}^{s+2|x|+O(1)}(x,y)\leq\operatorname{C}^{s}(x)+\operatorname{C}^{s}(y|x)+O(\log\operatorname{C}^{s}(x)).$
This is proven by simply running the program for $x$ and for $y$ given $x$. Also, the other inequality holds.
Theorem 1.2 ([5])
For all strings $x,y$ and number $s$,
$\operatorname{C}^{s^{\prime}}(x)+\operatorname{C}^{s^{\prime}}(y|x)\leq\operatorname{C}^{s}(x,y)+O(\log\operatorname{C}^{s}(x,y)),$
where $s^{\prime}=s+O(|xy|)$.
In the limit for large $s$, this inequality reduces to additivity for unbounded complexity. A weaker result from [14, 15] is well known: if $s^{\prime}=2s+O(|xy|+\log s)$, then $\operatorname{C}^{s^{\prime}}(x)+\operatorname{C}^{s^{\prime}}(y|x)\leq\operatorname{C}^{s}(x,y)+O(\log s+\log\operatorname{C}^{s}(x,y))$. Because of the $O(\log s)$ in the right side, the result (*1) for unbounded complexity is not a corollary. The proof of this version follows the one for unbounded complexity.222 There is a small caveat: the decompresser needs to iterate over pairs $(x,y)$ with $\operatorname{C}^{s}(x,y)<k$, and hence, iterate through all halting programs that use at most $s$ bits of memory. To check whether a program halts, one needs an additional $s$ bits of memory. The main idea of the proof of theorem 1.2 is to do this more efficiently using Sipser’s trick from [18]: given a halting configuration of a machine, consider all computations that terminate in this state, and organize all states in a tree. Then do a DFS search starting in this halting configuration.
The main result of this paper is that additivity of online information does hold for space-bounded complexity. This is indeed possible, because space bounded complexity is computable (in small space), see [5, lemma 3] or proposition 4 below. Moreover, the gap $s^{\prime}-s$ is very small.
Theorem 1.3
For all $x$ and $s$,
$\operatorname{C}_{\textnormal{ev}}^{s^{\prime}}(x)+\operatorname{C}_{\textnormal{odd}}^{s^{\prime}}(x)\leq\operatorname{C}^{s}(x)+O\big{(}\log s+\log|x|\big{)},$
where $s^{\prime}=s+6|x|+O(1)$.
Remark. It is not possible to drop the term $O(\log s)$, because otherwise, the limit for large $s$ would contradict theorem 1.1. Also, the proof gives a better precision of the big-O term at the right: $2d+4\log d+O(1)$ where $d=\operatorname{C}^{s}(s,\operatorname{C}^{s}(x),|x|)$.333 We believe it is possible to drop the term $O(d)$ when using prefix or monotone complexity, and still $s^{\prime}-s\leq O(|x|)$. However, the proof seems significantly harder. We hope to publish it in an extended version of this paper.
Proof structure. In the next section we define odd and even online semimeasures. We fix such semimeasures $\operatorname{P_{\textnormal{ev}}}$ and $\operatorname{P_{\textnormal{odd}}}$ that can be computed with $s+O(|x|)$ space and satisfy
$\operatorname{C}^{s}(x)\geq\log\frac{1}{\operatorname{P_{\textnormal{ev}}}(x)}+\log\frac{1}{\operatorname{P_{\textnormal{odd}}}(x)}.$
Let $k=\operatorname{C}^{s}(x)$. We prove a space-bounded version of the online coding theorem from [7]. This implies
$\operatorname{C}_{\textnormal{odd}}^{s}(x)$ $\leq\log\frac{1}{\operatorname{P_{\textnormal{odd}}}(x)}+O(\operatorname{C}^{s}(s,|x|,k))$
$\operatorname{C}_{\textnormal{ev}}^{s}(x)$ $\leq\log\frac{1}{\operatorname{P_{\textnormal{ev}}}(x)}+O(\operatorname{C}^{s}(s,|x|,k)).$
The theorem follows by combining the 3 inequalities.
2 Definitions
For background in Kolmogorov complexity, see [13, 17]. Typically, space-bounds in Kolmogorov complexity are defined and used up to arbitrary constant factors, see for example [9, 14, 15]. In [5], a definition is given with additive precision $O(1)$. We use it here as well.
We consider Turing machines that contain the following.
  • For each input, there is a separate one-directional read-only tape with end markers. (Both for the program and each auxiliary string in the definition of complexity.)
  • A one-directional write-only output tape.
  • Two binary stacks, with push, pop and isempty requests.
Remarks.
  • A computation uses space $s$ if the sum of the sizes of the 2 stacks is always at most $s$.
  • One may think of the 2 stacks as a standard work tape, where 1 stack represents the part left of the computation head and other stack the right part. Note that to make this analogy work, one needs to think about a binary work tape with some “end marker” that can be moved.444 Popping an element from the stack corresponds to shifting all elements on one side of the tape, and this operation can be done if an endmarker is available. This definition allows for cleaner statements and proofs of space-bounded additivity (with $s^{\prime}=s+O(|xy|)$ without extra $O(\log s)$ term).
  • In proofs, one may use a more general type of machine, which is equipped with several standard binary work tapes that have end markers. Consider a machine with $k$ such tapes. Also, consider a computation on which at most $t_{1},\ldots,t_{k}$ bits of these tapes are used, and the total stack size is at most $s$. This computation is simulated on a machine of the above type using space $s+\sum_{i\leq k}t_{i}+2\lceil\log t_{i}\rceil$.
Definition 1
Let $x_{1},y_{1},\ldots,x_{k},y_{k}$ be strings. The Kolmogorov complexity of tasks $x_{1}{\leftarrow}y_{1}$; …; $x_{k}{\leftarrow}y_{k}$ with space bound $s$ relative to machine $V$ is
$\operatorname{C}^{s}_{V}(x_{1}{\leftarrow}y_{1};\ldots;x_{k}{\leftarrow}y_{k})=\min\big{\{}|p|:\forall j\mathop{\leq}k\big{[}V(p,y_{j})=x_{j}\textnormal{ using space }s\big{]}\big{\}}.$
Proposition 1
There exists a machine $U$ with the following property. For each machine $V$ there exists a constant $c$ such that for all $s,k$ and strings $x_{1},y_{1},\ldots,x_{k},y_{k}$,
$\operatorname{C}_{U}^{s+c}(x_{1}{\leftarrow}y_{1};\ldots;x_{k}{\leftarrow}y_{k})\leq\operatorname{C}_{V}^{s}(x_{1}{\leftarrow}y_{1};\ldots;x_{k}{\leftarrow}y_{k})+c.$
The proof follows the one of [5, proposition 1], where a universal machine is constructed in which a self-delimiting description of the simulated machine $V$ is prepended to its program. This description is stored on an auxiliary tape. The constant $c$ is the extra space needed by the auxiliary tape (in the simulation on the stacks).
We fix such a machine $U$ and drop subscripts $U$ in all complexities.
Definition 2
Let $x,y\in\{0,1\}^{*}$. The conditional complexity is $\operatorname{C}^{s}(x|y)=\operatorname{C}^{s}(x{\leftarrow}y)$. If $y$ is empty, we write $\operatorname{C}^{s}(x)$. We use the prefix code $\bar{y}=0y_{1}\cdots 0y_{|y|}1$ and encode a tuple of strings $(t_{1},\ldots,t_{e})$ as $\bar{t}_{1}\bar{t}_{2}\cdots\bar{t}_{e-1}t_{e}$. The conditional complexity with tuples is defined using this encoding, for example $\operatorname{C}^{s}(y,z|x)=\operatorname{C}^{s}(\bar{y}z|x)$.
The even complexity of $x$ is
$\operatorname{C}_{\textnormal{ev}}^{s}(x)$ $=\min\big{\{}|p|:\forall j\leq|x|/2\big{[}U(p,x_{1}x_{3}\cdots x_{2j-1})=x_{2j}\textnormal{ using space }s\big{]}\big{\}}$
$=\operatorname{C}^{s}(x_{2}{\leftarrow}x_{1};\,x_{4}{\leftarrow}x_{1}x_{3};\,\ldots).$
The odd complexity of $x$ is defined similarly. For unbounded complexity, thus $s=+\infty$, we drop $s$, for example $\operatorname{C}^{\infty}(x|y)=\operatorname{C}(x|y)$.
Remark. The decision variant of the Kolmogorov complexity of $x$, is the minimal length of a program that prints some string or infinite sequence that has $x$ as prefix. It would be fair to compare $\operatorname{C}_{\textnormal{ev}}(x)+\operatorname{C}_{\textnormal{odd}}(x)$ to this variant, since the odd and even complexity do not require to indicate where $x$ stops. However, the decision and plain complexities differ by at most $O(\log\operatorname{C}(|x|))$, and this is less than the precision $O(\log(s|x|))$ of the main result. Hence, there is no reason to deviate from the more standard notions.
3 Even and odd online measures
Definition 3
Let $P$ be a mapping from strings to the real interval $[0,1]$.
  • $P$ is a semimeasure if for all $x$,
    $P(x)\geq P(x0)+P(x1).$
  • $P$ is an even semimeasure if for all $x$,
    if $|x|$ is even $\Rightarrow$ $P(x)=P(x0)=P(x1)$
    if $|x|$ is odd $\Rightarrow$ $P(x)\geq P(x0)+P(x1).$
  • $P$ is an odd semimeasure if for all $x$,
    if $|x|$ is odd $\Rightarrow$ $P(x)=P(x0)=P(x1)$
    if $|x|$ is even $\Rightarrow$ $P(x)\geq P(x0)+P(x1).$
An online coding theorem for lower-semicomputable even semimeasures was proven in [7]. However, it has a logarithmic term in the right side. Below, we prove a variant for computable even semimeasures without this term.
Proposition 2
For each computable even semimeasure $\operatorname{P_{\textnormal{ev}}}$ there exists a constant $c$ such that for all $x$,
$\operatorname{C}_{\textnormal{ev}}(x)\leq\log\frac{1}{\operatorname{P_{\textnormal{ev}}}(x)}+c.$
The goal of this section is to prove a version of this theorem with a space-bound. We use the following.
Definition 4
The complexity $\operatorname{C}^{\tilde{s}}(F)$ of $F:\{0,1\}^{*}\rightarrow\mathbb{R}$ with space bound $\tilde{s}:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$ is the minimal length of a program on $U$ that on input $(k,x)$ uses space $\tilde{s}(k,|x|)$ and outputs an integer $N$ such that $|F(x)-N2^{-k}|\leq 2^{-k}$, where $k$ is given in binary. If no such program exists, then $\operatorname{C}^{\tilde{s}}(F)=+\infty$.
Proposition 3
Assume $\tilde{s}$ is non-decreasing in both arguments. There exists a $c$ such that for every even semimeasure $\operatorname{P_{\textnormal{ev}}}$ and string $x$,
$\operatorname{C}_{\textnormal{ev}}^{s^{\prime}}(x)\leq\log\frac{1}{\operatorname{P_{\textnormal{ev}}}(x)}+O(\operatorname{C}^{\tilde{s}}(\operatorname{P_{\textnormal{ev}}})),$
where $s^{\prime}=\tilde{s}(\big{\lceil}\log\frac{8|x|}{\operatorname{P_{\textnormal{ev}}}(x)}\big{\rceil},|x|)+|x|+O(\log|x|)$.
A similar result for $\operatorname{P_{\textnormal{odd}}}$ holds with a similar proof.
Proof (of proposition 2)
We first explain that any real even semimeasure can be transformed to a rational one that is at most a constant factor smaller. Indeed, we may assume that $\operatorname{P_{\textnormal{ev}}}(x)\geq 2^{-\lfloor|x|/2\rfloor-1}$, otherwise, we replace it by $2^{-\lfloor|x|/2\rfloor-1}+\tfrac{1}{2}\operatorname{P_{\textnormal{ev}}}(x)$. Then, it is sufficient to round the value down with precision $2^{-|x|}$ for even $|x|$, and rescale recursively. This decreases the semimeasure by at most a factor $\prod_{i=1}^{\infty}(1-2^{-i})>0.$
We now assume that $\operatorname{P_{\textnormal{ev}}}$ is rational. Let the cumulative semimeasure $\operatorname{PC}(y)$ represent the measure of all $z$ with $|z|=|y|$ such that $0.z<0.y$, more precisely, let PC(empty $\operatorname{PC}(\textnormal{empty str})=0$ and
$\operatorname{PC}(yb)=\begin{cases}\operatorname{PC}(y)&\text{ if }|y|\textnormal{ is even or }b=0,\\ \operatorname{PC}(y)+\operatorname{P_{\textnormal{ev}}}(y0)&\textnormal{ otherwise, thus $|y|$ is odd and $b=1$.}\end{cases}$
The decompresser depends on this function. For notational convenience, consider a decompresser in the definition of even complexity that also takes the previous answers as input. For example, in $\operatorname{C}_{\textnormal{ev}}(y)$, on input $y_{1}y_{2}y_{3}y_{4}y_{5}$ the program outputs $y_{6}$, (instead of input $y_{1}y_{3}y_{5}$).
Machine $V$ on input $p$ and $y$: if $2^{-|p|}p<\operatorname{PC}(y1)$ output 0, otherwise output 1.
Recall that $\operatorname{P_{\textnormal{ev}}}$ is a computable rational function. Hence, also $\operatorname{PC}$ is a computable rational function, because $\operatorname{PC}(y)$ equals the sum of at most $|y|/2$ values $\operatorname{P_{\textnormal{ev}}}(\cdot)$. Thus $V$ computes 2 rational numbers and outputs the result of their comparison.
$b=0$$\operatorname{PC}(y)$$\operatorname{PC}(y1)$$\operatorname{P_{\textnormal{ev}}}(y)$$\operatorname{P_{\textnormal{ev}}}(y0)$$\operatorname{P_{\textnormal{ev}}}(y1)$$p_{x}$
We verify that if $y$ is a strict prefix of $x$ with odd length and if
$p2^{-|p|}\in\big{[}\operatorname{PC}(x),\operatorname{PC}(x)+\operatorname{P_{\textnormal{ev}}}(x)\big{)},$
then $V(p,y)=x_{|y|+1}$. This is quite logical, but we give the details anyway. Fix a prefix $z$ of $x$, and let $b=x_{|z|+1}$. If $b=0$, then
$\big{[}\operatorname{PC}(zb),\operatorname{PC}(zb)+\operatorname{P_{\textnormal{ev}}}(zb)\big{)}\;\subseteq\;\big{[}\operatorname{PC}(z),\operatorname{PC}(z)+\operatorname{P_{\textnormal{ev}}}(z)\big{)}.$
This also holds for $b=1$, by expanding the definition of $\operatorname{PC}(z1)$. It also holds for $|z|$ of even length. For $b=x_{|y|+1}$, it holds that
$p2^{-|p|}\;\in\;\big{[}\operatorname{PC}(x),\operatorname{PC}(x)+\operatorname{P_{\textnormal{ev}}}(x)\big{)}\;\subseteq\;\big{[}\operatorname{PC}(yb),\operatorname{PC}(yb)+\operatorname{P_{\textnormal{ev}}}(yb)\big{)}.$
We verify that $b=0$ if and only if $2^{-|p|}p<\operatorname{PC}(y1)$:
– If $b=0$, then $p2^{-|p|}<\operatorname{PC}(y0)+\operatorname{P_{\textnormal{ev}}}(y1)=\operatorname{PC}(y1)$.
– If $b=1$, then $p2^{-|p|}\geq\operatorname{PC}(y1)$.
This implies $V(p,y)=b$.
To finish the proof, consider any $x$ with $\operatorname{P_{\textnormal{ev}}}(x)>0$. Let $k_{x}=\lceil\log\frac{1}{\operatorname{P_{\textnormal{ev}}}(x)}\rceil$ and let
$p_{x}=\lceil 2^{k_{x}}\cdot\operatorname{PC}(x)\rceil$
be a number in binary with $k_{x}$ bits (if needed, add leading zeros). Note that $2^{-k_{x}}p_{x}$ lies in the interval $\big{[}\operatorname{PC}(x),\operatorname{PC}(x)+\operatorname{P_{\textnormal{ev}}}(x)\big{)}$, because $\operatorname{P_{\textnormal{ev}}}(x)\geq 2^{-k_{x}}$. Thus,
$\operatorname{C}_{\textnormal{ev},V}(x)\leq|p_{x}|=k_{x}=\Big{\lceil}\log\frac{1}{\operatorname{P_{\textnormal{ev}}}(x)}\Big{\rceil}.$
Proof (of proposition 3)
Now $\operatorname{P_{\textnormal{ev}}}$ is a real valued function and computable with space bound $\tilde{s}$. The idea is to use the program $p^{\prime}_{x}=p_{x}01$, which represents a number in the interval $\big{[}\operatorname{PC}(x),\operatorname{PC}(x)+\operatorname{P_{\textnormal{ev}}}(x)\big{)}$ that is separated from the borders by at least $2^{-|p_{x}|-2}$. We use the same algorithm for $V$, but $\operatorname{PC}(y1)$ is calculated with precision $2^{-|p_{x}|-2}$.
For this, all of the at most $|y|/2$ used values of $\operatorname{P_{\textnormal{ev}}}$ are computed with $k_{x}+2+\lceil\log|y|\rceil$ bits precision, where $k_{x}=\lceil\log\operatorname{P_{\textnormal{ev}}}(x)\rceil$. This requires space
$\tilde{s}(k_{x}+2+\lceil\log|x|\rceil,|x|)+k_{x}+\log|x|+O(1).$
4 Proof of theorem 1.3
In [5, lemma 3] it is shown that on input $x$ and $s$, one can compute $\operatorname{C}^{s}(x)$ in space $s+O(|x|)$. With a similar proof we obtain the following.
Proposition 4
There exists a machine that on input $x,k,s$ decides whether $\operatorname{C}^{s}(x)<k$ using space $s+2|x|+O(\log|x|)$.
The proof of theorem 1.3 relies on defining an odd and an even semimeasure. They are derived from the following (standard) semimeasure. Fix $s,k$ and $n$. Let
$Q(y)=\frac{\big{|}\big{\{}z:|yz|=n\textnormal{ and }\operatorname{C}^{s}(yz)\mathop{<}k\big{\}}\big{|}}{2^{k}}.$
Remarks.
– There are at most $12^{k}-$ 1 strings with complexity less than $k$, thus $Q(\text{empty})<1$.
$Q$ is indeed a semimeasure with $Q(x)=0$ if $|x|>n$.
– If $\operatorname{C}^{s}(x)<k$, then $-\log Q(x)\leq k$.
– The space needed to evaluate $Q(y)$ exactly is
$(n+O(\log n))+(s+2n+O(\log n)),$
where the first term, is for storing the counter $z$ (in the definition of $Q$), and the second is the bound of proposition 4 for checking $\operatorname{C}^{s}(yz)<k$.
For $x$ of even length $\ell$, the odd and even semimeasures are
$\operatorname{P_{\textnormal{odd}}}(x)$ $=Q(x_{1})\cdot\frac{Q(x_{1}x_{2}x_{3})}{Q(x_{1}x_{2})}\cdot\ldots\cdot\frac{Q(x_{1}\cdots x_{\ell-2}x_{\ell-1})}{Q(x_{1}\cdots x_{\ell-2})}$
$\operatorname{P_{\textnormal{ev}}}(x)$ $=\frac{Q(x_{1}x_{2})}{Q(x_{1})}\cdot\frac{Q(x_{1}x_{2}x_{3}x_{4})}{Q(x_{1}x_{2}x_{3})}\cdot\ldots\cdot\frac{Q(x_{1}\cdots x_{\ell-1}x_{\ell})}{Q(x_{1}\cdots x_{\ell-1})},$
and these semimeasures are then uniquely defined on strings of odd lengths.
Remarks.
$Q(x)=\operatorname{P_{\textnormal{odd}}}(x)\operatorname{P_{\textnormal{ev}}}(x)$.
– For $\tilde{s}(k^{\prime},n^{\prime})=s+3n^{\prime}+k^{\prime}+O(\log n^{\prime})$, we prove that
$\operatorname{C}^{\tilde{s}}(\operatorname{P_{\textnormal{odd}}})=\operatorname{C}^{s}(s,n,k)+O(1).$
Indeed, if $|y|>n$, then $\operatorname{P_{\textnormal{odd}}}(y)=0$. Otherwise, $\operatorname{P_{\textnormal{odd}}}$ is the product of $|y|=n^{\prime}$ numbers, and each is computed and stored with precision $k^{\prime}+\lceil\log n^{\prime}\rceil$.
Now we finish the proof of the main theorem. For some $x$ and $s$, consider $Q$ with parameters $s=s,\,k=\operatorname{C}^{s}(x)+1$ and $n=|x|$. We may assume $k\leq n$, otherwise the theorem is trivial. For $s^{\prime}=\tilde{s}(n+3+\lceil\log n\rceil,n)+n+O(\log n)$,
$\operatorname{C}^{s}(x)+1\geq-\log Q(x)=-\log\Big{(}\operatorname{P_{\textnormal{odd}}}(x)\,\operatorname{P_{\textnormal{ev}}}(x)\Big{)}\geq\operatorname{C}_{\textnormal{ev}}^{s^{\prime}}(x)+\operatorname{C}_{\textnormal{odd}}^{s^{\prime}}-O(\operatorname{C}^{s}(s,n,k)).$
Since $s^{\prime}\leq s+6n+O(1)$, this finishes the proof of theorem 1.3.